a cylindrical vessel, without lid, has to be tin-coated on its both sides. if the radius of the base is 70cm and its height is 1.4m, calculate the cost of tin coating at the rate of Rs3.50 per 1000m2.

Answer by gonzo(654)

assuming no thickness to the sides of the can, i would estimate the answer to be as follows:
h = 1.4*m = height of the can
r = 70*cm = 7000*m = radius of the can (1 centimeter = 100 meters)
surface area on both sides of the can would be 2 * circumference of the base * the height.
c = circumference of the base = circumference of a circle = 2*pi*r.
s = surface area on both sides of the can.
s = 2*c*h = 2*2*pi*r*h = 2*pi*7000*m*1.4*m = 2*pi*9800*m^2 = pi*19600*m^2 = 61575.21601*m^2
-----
not sure what Rs3.50 per 1000m2 means but i think it means:
cost of tin coating = $3.50 per 1000 square meters of surface area.
ctc = cost of tin coating.
ctc = 61575.21601*m^2 * $3.5 / (1000*m^2) = $215.513256 = $215.51
-----